OpenStudy (anonymous):

ennifer has been saving for college for 57 months. The first month, she saved $11. She was able to save more money each month than the month before. She ended up saving $19,779.00. How much more did she save each month?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

What do you think? :)

OpenStudy (anonymous):

I'm thinking the final amount is the sum of an arithmetic series and they want you to find the common difference of the sequence

OpenStudy (anonymous):

I agree!

OpenStudy (anonymous):

But I just dont know how to do it hahah

OpenStudy (anonymous):

Use this one to find \(a_{57}\), the amount saved in the 57th month. \[S_{57}=\frac{ n }{ 2 }(a_1+a_{57})\] n = 57 and \(a_1=11\)

OpenStudy (anonymous):

ok so s57=28.5(11+a57)

OpenStudy (anonymous):

Right and S_57 = 19779

OpenStudy (anonymous):

now what do we do?

OpenStudy (anonymous):

solve for \(a_{57}\)

OpenStudy (anonymous):

What would I divide by?

OpenStudy (anonymous):

28.5

OpenStudy (anonymous):

694

OpenStudy (anonymous):

subtract 11

OpenStudy (anonymous):

683

OpenStudy (anonymous):

that's the 57th term. This is the formula to find d\[a_n=a_1+d(n-1)\] \[683=11+d(57-1)\]

OpenStudy (anonymous):

683=57d-d+11

OpenStudy (anonymous):

11.78=answer

OpenStudy (anonymous):

But the answer choices are 11.00 11.50 12.00 12.50

OpenStudy (anonymous):

683=11 + 56d

OpenStudy (anonymous):

672 = 56d

OpenStudy (anonymous):

12:)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks :) If I open anothe question, will u answer it?